### v1

@@ Abstract
The resolving ability of wide-field fluorescence microscopy is fundamentally limited by out-of-focus background owing to its low axial resolution, particularly for densely labeled biological samples. To address this, we developed ET2dNet, a deep learning-based EPI-TIRF cross-modality network that achieves TIRF-comparable background subtraction and axial super-resolution from a single wide-field image without requiring hardware modifications. The model employs a physics-informed hybrid architecture, synergizing supervised learning with registered EPI-TIRF image pairs and self-supervised physical modeling via convolution with the point spread function. This framework ensures exceptional generalization across microscope objectives, enabling few-shot adaptation to new imaging setups. Rigorous validation on cellular and tissue samples confirms ET2dNet’s superiority in background suppression and axial resolution enhancement, while maintaining compatibility with deconvolution techniques for lateral resolution improvement. Furthermore, by extending this paradigm through knowledge distillation, we developed ET3dNet, a dedicated three-dimensional reconstruction network that produces artifact-reduced volumetric results. ET3dNet effectively removes out-of-focus background signals even when the input image stack lacks the source of background. This framework makes axial super-resolution imaging more accessible by providing an easy-to-deploy algorithm that avoids additional hardware costs and complexity, showing great potential for live cell studies and clinical histopathology.
@@ 1 Introduction
Fluorescence microscopy serves as a cornerstone technology in cell biology, enabling visualization of tissue and subcellular structures. However, wide-field (WF) fluorescence microscopy suffers from substantial out-of-focus background and limited axial resolution[stephens_light_2003, liu_breaking_2018], which degrade performance on deep cell and tissue imaging. To address these limitations, both hardware-based microscopy enhancements and computational algorithms were developed.
@@
Inspired by deep learning cross-modality frameworks, we developed an Epifluorescence (EPI)-TIRF cross-modality deep neural network capable of suppressing out-of-focus background and enhancing axial-resolution from a single WF image, achieving a TIRF-comparable imaging quality. Crucially, it overcame TIRF’s shallow imaging depth limitation, enabling deep cell and tissue imaging. We enhanced generalization through few-shot learning by integrating image degradation modeling into the network architecture, which was validated against TIRF and confocal GT. Furthermore, we employed knowledge distillation [wang_knowledge_2022] to train a 3D reconstruction network requiring only 6 optical slices, ensuring rapid and faithful volumetric reconstruction. Our work provided a powerful and user-friendly new tool for enhancing WF microscopy imaging, significantly facilitating biological and medical research. It simultaneously advanced fundamental methodologies for deep learning-based super-resolution microscopy.
@@ 2 Results
2.1 EPI-TIRF cross-modality enables optical sectioning and axial super-resolution for a single WF image
Figure 1: EPI-TIRF cross-modality network for a single WF image. a EPI and TIRF illumination modes. b The architecture of ET2dNet. c Representative F-actin labeled C6 cell image of EPI(first column), Sparse-Deconvolution reconstruction (second column), Dark Sectioning reconstruction (third column), ET2dNet fake TIRF (fourth column), Sparse-Deconvolution reconstruction of ET2dNet fake TIRF (fifth column), and TIRF (sixth column). The magnified view of region in white box is displayed at the bottom. d EPI z-scan stack of white dashed box highlighted region in (c), with 200nm step size, and corresponding reconstruction. e Average MS-SSIM (to TIRF) of processed images in test dataset, by either Sparse-Deconvolutionm, Dark Sectioning and ET2dNet. f Representative images of human brain tissue section stained with an antibody against GFAP (provided by ref.[wernersson_deconwolf_2024]), EPI(first column), Sparse-Deconvolution reconstruction (second column), Dark Sectioning reconstruction (third column), ET2dNet fake TIRF (fourth column), Sparse-Deconvolution reconstruction of ET2dNet fake TIRF (fifth column), and confocal (sixth column). The magnified view of region in white box is displayed at the bottom, with marker line and corresponding intensity profile. Scale bar: 5 $\rm\mu m$.
@@ 2.1 EPI-TIRF cross-modality enables optical sectioning and axial super-resolution for a single WF image
We developed the network, denoted ET2dNet, with a hybrid architecture. ET2dNet comprised dual branches for self-supervised and supervised learning to generate two outputs, fake GT and background, as shown in Fig.1 b. The in-focus fake GT was convolved with the point spread function (PSF) to generate fake TIRF. A supervised loss was computed by comparing fake and real TIRF (see Method 4.6). Then the fake TIRF and background were combined to generate a fake EPI, enabling self-supervised training through comparison with real EPI. For the network backbone, we implemented EViT-UNet [li_evit-unet_2025], a state-of-the-art (SOTA) segmentation network.
The hybrid network architecture drew conceptual inspiration from the classical deconvolution theory[bertero_introduction_2021], where iterative gradient descent minimizes reconstruction errors through regularization-guided optimization by incorporating a prior. In this deep learning framework, the supervised learning branch encoded biological priors through training on real EPI-TIRF pairs, and the self-supervised learning branch explicitly modeled the image degradation process via PSF convolution. The generalization capability of the architecture, enhanced by its self-supervised component (as discussed in the latter section), aligns with recent advances in microscopic[qiao_zero-shot_2024] and natural[chen_low-res_2024] image super-resolution.
@@
2.1.2 ET2dNet predictions show efficient background subtraction and axial-resolution enhancement
@@
ET2dNet removed background and achieved axial super-resolution. We implemented ET2dNet on the EPI image and compared the output (fake TIRF) with the real TIRF image, as shown in Fig.1 c. Visually, the fake TIRF showed minimal background and high contrast. Then we focused on the region highlighted by the white dashed box, where the EPI and TIRF images revealed distinct characteristics of the microfilament structure. EPI image showed a continuous microfilament, while the TIRF image depicts a discontinuous structure. Given the shallow penetration depth of TIRF illumination (approximately 50-200 nm) at the bottom in the z-direction, we inferred that the microfilament formed an arch-like structure, with the middle part positioned away from the bottom. This structure served as a typical example to assess the axial-resolution enhancement achieved by ET2dNet. We then acquired six EPI images along the z-axis, spaced 200 nm apart, as illustrated in Fig.1 d. Visually, these images appeared nearly indistinguishable. However, the corresponding ET2dNet outputs exhibited distinct differences and collectively reconstruct an arch-like structure, demonstrating the enhanced resolution achieved by ET2dNet. The corresponding full image stacks are provided in Supplementary Information Fig. S1.
@@
We compared the performance among ET2dNet, SOTA deconvolution (Sparse Deconvolution[zhao_sparse_2022]) and background removal (Dark Sectioning[cao_dark-based_2025]) methods, as shown in Fig.1 c. The Sparse Deconvolution gave a sharper reconstruction, while the out-of-focus signal still existed. The Dark Sectioning could efficiently remove the background, but barely enhanced the axial resolution, which was clearly shown by the region of the arch-like structure. It should also be noted that, ET2dNet was not specifically designed to implement deconvolution, but it was compatible with deconvolution methods, as shown in the sixth row in Fig.1 d. By directly implementing Sparse Deconvolution on the fake TIRF of ET2dNet, both lateral and axial resolution were enhanced. We further examined ET2dNet on a test dataset containing 20 images, generating a better average SSIM (compared to Sparse Deconvolution and Dark Sectioning), as shown in Fig.1 e. Additional examples of F-actin labeled U87 cells are shown in Extended Data Fig.1. We also compared other deconvolution and background removal methods, including RL deconvolution[richardson_bayesian-based_1972, lucy_iterative_1974], Wavelet-based background and noise subtraction (WBNS)[hupfel_wavelet-based_2021], Background Filtering (BF)[mo_quantitative_2023], while ET2dNet outperformed all these methods, see Extended Data Fig.2. Brief introductions and implementation details of the above algorithms are provided in Supplementary Information Note 1. The details about the combination of ET2dNet and Sparse deconvolution are provided in Supplementary Information Note 2.
@@
The performance of ET2dNet was further assessed using WF-confocal image pairs of human brain tissue sections stained with an antibody against GFAP (provided by ref.[wernersson_deconwolf_2024]), as shown in Fig.1 f. Note that the WF image was collected using 60X 1.4NA oil objective (with pixel size 107.9nm), which was slightly different from the training dataset using 60X 1.5NA oil objective (with pixel size 97.5nm). In such case, ET2dNet could still efficiently remove background, outperforming the Sparse Deconvolution method and Dark Sectioning. As indicated by the region in the white box and the corresponding intensity profile, Sparse Deconvolution failed to suppress background, while Dark Sectioning erroneously eliminated signals. In contrast, ET2dNet achieved a background-free result while preserving critical structural information, demonstrating high consistency with confocal GT.
2.2 Generalization of ET2dNet and few-shot learning
Figure 2: Generalization and finetuning of ET2dNet. a Representative F-actin labeled C6 cell image of EPI(first column), ET2dNet 60X deconvolution reconstruction (second column), finetuned ET2dNet 100X deconvolution reconstruction (third column), and TIRF (fourth column). The magnified view of region in white box is displayed at the bottom. b Statistical comparisons of ET2dNet 60X/100X in terms of SSIM and PSNR across 20 EPI-TIRF image pairs. c Outputs of base and hybrid models at different fine-tuning epochs (0, 5, 10, 15), for the region in white box in (a). d Statistical comparisons of base (black) and hybrid (red) models in terms of MS-SSIM during the finetune(mean ± $\sigma$ confidence interval, smoothed over 3 epochs; 20 image pairs per model scale). e Representative images of human brain tissue section stained with an antibody against GFAP of EPI(first column), ET2dNet 60X output(second column), finetuned ET2dNet 100X output(third column), confocal(forth column), focus at 0 $\mu m$ (top) and 0.75 $\mu m$ (bottom). f The normalized intensity profile along the marker lines corresponding to (d). Scale bar: 6 $\rm\mu m$ in (e) and top of (a), 3 $\rm\mu m$ in (c) and bottom of (a).
@@ 2.2 Generalization of ET2dNet and few-shot learning
Our deep learning method aimed to generalize across WF microscopes featuring diverse objectives (e.g., varying numerical aperture and magnification). We hypothesized ET2dNet exhibited robust performance under similar PSFs (though non-identical), as validated in Fig.1f. Here, input images acquired using a 60X/1.4NA objective were processed by a network trained on datasets acquired using a 60X/1.5NA objective. For systems exhibiting substantially different PSFs, fine-tuning was required. Implementation of ET2dNet on WF images acquired with a 100X/1.5NA objective revealed effective background removal but negligible enhancement in axial-resolution, as shown in Fig.2a (denoted ET2dNet 60X). Specifically, at an arch-like structure (white dashed box), continuous microfilaments in the ET2dNet 60X output indicated a lack of axial-resolution improvement.
@@
To improve the axial-resolution enhancement capability of ET2dNet, we finetuned it on a small dataset of 20 paired EPI-TIRF images. The finetuning process achieved rapid convergence, yielding ET2dNet 100X, which significantly outperformed ET2dNet 60X at axial-resolution enhancement. For instance, at the arch-like structure indicated by the white box in Fig.2 a, the ET2dNet 100X predicted discontinuous microfilaments, aligning with the TIRF GT. Quantitative validation across 20 test image pairs further confirmed improved PSNR and MS-SSIM after fine-tuning (Fig.2b).
@@
Furthermore, we demonstrated that the hybrid network architecture enhances generalization and advances the finetuning. An ablation study investigating the impact of the self-supervised learning branch was presented in Fig.2 c and d. Both the base model (without self-supervised learning branch) and the hybrid model were trained on the 60X dataset and subsequently fine-tuned on the 100X dataset. Visually, the hybrid model regained its axial-resolution enhancement capability within 5 epochs during fine-tuning, whereas the base model required more than 20 epochs, as shown in Fig. 2c (the full images are provided in Supplementary Information Fig. S2). Statistical comparison in Fig. 2d revealed the hybrid model initially achieved higher average MS-SSIM (to TIRF GT in the test dataset) and converged faster to a superior final value, reinforcing our conclusion. Ablation experiments conducted on networks with different model sizes (represented by S2, S1, S0 in Fig. 2d) consistently corroborated the above findings, underscoring the hybrid model’s strong transferability and superior generalization capacity.
@@
Axial resolution enhancement was further validated on images of stained human brain tissue sections (Fig. 2e). Comparative analysis of ET2dNet implementations (60X vs. 100X) on z-stack (containing two 0.75 µm axial spacing slices) revealed critical performance differences. While the 60X model provided partial background suppression, the fine-tuned 100X model substantially eliminated out-of-focus signals. Notably, two microfilaments (indicated by green and blue arrows) located at the 0 µm and 0.75 µm planes exhibited cross-plane interference in EPI images. As evidenced by intensity profiles (Fig. 2e), both confocal microscopy and ET2dNet 60X failed to resolve these structures, whereas ET2dNet 100X successfully isolated the overlapping signals. This demonstrated ET2dNet’s capability to surpass conventional confocal microscopy in axial-resolution, delivering effective optical sectioning below the diffraction limit.
2.3 Artifact-reduced 3D reconstruction network

### v2

@@ Abstract
The resolving ability of widefield fluorescence microscopy is fundamentally limited by out-of-focus background owing to its low axial resolution, particularly for densely labeled biological samples. Although total internal reflection fluorescence (TIRF) microscopy provides strong near-surface sectioning, they are intrinsically restricted to shallow imaging depths. Here we present computational TIRF (cTIRF), a deep learning-based imaging modality that generates TIRF-like sectioned images directly from conventional widefield epifluorescence measurements without any optical modification. By integrating a physics-informed forward model into network training, cTIRF achieves effective background suppression and axial resolution enhancement while maintaining consistency with the measured widefield data. We demonstrate that cTIRF recovers near-surface structures with performance comparable to experimental TIRF, and further enables both single-frame and volumetric sectioned reconstruction in densely labeled samples where conventional TIRF fails. This work establishes cTIRF as a practical and deployable alternative to hardware-based optical sectioning in fluorescence microscopy, enabled by rapid adaptation to new imaging systems with minimal calibration data.
@@ 1 Introduction
Fluorescence microscopy serves as a cornerstone technology in cell biology, enabling visualization of tissue and subcellular structures. However, widefield (WF) fluorescence microscopy suffers from substantial out-of-focus background and limited axial resolution[stephens_light_2003, liu_breaking_2018], which degrade performance on deep cell and tissue imaging. To address these limitations, both hardware-based microscopy enhancements and computational algorithms were developed.
@@
Inspired by cross-modality learning concepts, we introduce computational TIRF (cTIRF) as a computational imaging modality that capable of suppressing out-of-focus background and enhancing axial-resolution from a single widefield epifluorescence image. cTIRF achieved a TIRF-comparable imaging quality, while overcoming the intrinsic depth limitation of hardware-based TIRF and enabling imaging throughout whole cells and tissue sections. To facilitate practical deployment, we enable rapid adaptation of cTIRF to new imaging systems through few-shot learning, achieved by integrating physically motivated image degradation modeling into the reconstruction framework. The cTIRF representations are quantitatively validated against experimentally measured TIRF and confocal microscopy. Furthermore, we extend cTIRF to volumetric reconstruction by employing knowledge distillation [wang_knowledge_2022] to train a three-dimensional network that requires only six optical sections, enabling fast and faithful three-dimensional imaging under limited axial sampling. Together, this work establishes cTIRF as a powerful and user-friendly computational imaging modality for enhancing widefield fluorescence microscopy, with broad applicability to biological and biomedical imaging.
@@ 2 Results
2.1 Computational TIRF enables optical sectioning and axial super-resolution for a single WF image
Figure 1: Overview of computational TIRF for a single WF image. a EPI and TIRF illumination modes. b The architecture of ET2dNet. c Representative F-actin labeled C6 cell image of EPI(first column), Sparse-Deconvolution reconstruction (second column), Dark Sectioning reconstruction (third column), cTIRF (fourth column), Sparse-Deconvolution reconstruction of cTIRF (fifth column), and experimental measured TIRF (sixth column). The magnified view of region in white box is displayed at the bottom. d EPI z-scan stack of white dashed box highlighted region in (c), with 200nm step size, and corresponding reconstruction. e Average MS-SSIM (to TIRF) of processed images in test dataset, by either Sparse-Deconvolution, Dark Sectioning and cTIRF. f Representative images of human brain tissue section stained with an antibody against GFAP, EPI(first column), Sparse-Deconvolution reconstruction (second column), Dark Sectioning reconstruction (third column), cTIRF (fourth column), Sparse-Deconvolution reconstruction of cTIRF (fifth column), and confocal (sixth column). The magnified view of region in white box is displayed at the bottom, with marker line and corresponding intensity profile. Scale bar: 5 $\rm\mu m$.
@@ 2.1 Computational TIRF enables optical sectioning and axial super-resolution for a single WF image
We developed the EPI-TIRF cross-modality network, denoted ET2dNet, with a hybrid architecture. ET2dNet comprised dual branches for self-supervised and supervised learning to generate two outputs, pesudo-GT and background, as shown in Fig.1 b. The in-focus pesudo-GT was convolved with the point spread function (PSF) to generate computational TIRF. For simplification, we define pesudo-GT and computational TIRF (cTIRF) as the output produced by our model; unless otherwise specified, pesudo-GT and cTIRF images are generated using ET2dNet (2D) or ET3dNet (3D reconstruction network, see latter sections). A supervised loss was computed by comparing cTIRF and experimental measured TIRF. Then the cTIRF and background were combined to generate a reconstructed EPI, enabling self-supervised training through comparison with experimental measured EPI. For the network backbone, we implemented EViT-UNet [li_evit-unet_2025], a state-of-the-art (SOTA) segmentation network. Details are shown in Methods 4.4.
The hybrid network architecture drew conceptual inspiration from the classical deconvolution theory[bertero_introduction_2021], where iterative gradient descent minimizes reconstruction errors through regularization-guided optimization by incorporating a prior. In this deep learning framework, the supervised learning branch encoded biological priors through training on experimental measured EPI-TIRF pairs, and the self-supervised learning branch explicitly modeled the image degradation process via PSF convolution. The generalization capability of the architecture, enhanced by its self-supervised component (as discussed in the latter section), aligns with recent advances in microscopic[qiao_zero-shot_2024] and natural[chen_low-res_2024] image super-resolution.
@@
2.1.2 cTIRF shows efficient background subtraction and axial-resolution enhancement
@@
cTIRF removed background and achieved axial super-resolution. We implemented ET2dNet on the EPI image and compared the output (cTIRF) with the experimental measured TIRF image, as shown in Fig.1 c. Visually, the cTIRF showed minimal background and high contrast. Then we focused on the region highlighted by the white dashed box, where the measured EPI and TIRF images revealed distinct characteristics of the microfilament structure. EPI image showed a continuous microfilament, while the TIRF image depicts a discontinuous structure. Given the shallow penetration depth of TIRF illumination (approximately 50-200 nm) at the bottom in the z-direction, we inferred that the microfilament formed an arch-like structure, with the middle part positioned away from the bottom. This structure served as a typical example to assess the axial-resolution enhancement achieved by cTIRF. We then acquired six EPI images along the z-axis, spaced 200 nm apart, as illustrated in Fig.1 d. Visually, these images appeared nearly indistinguishable. However, the corresponding cTIRFs exhibited distinct differences and collectively reconstruct an arch-like structure, demonstrating the enhanced resolution achieved by cTIRF. The corresponding full image stacks are provided in Supplementary Information Fig. S1.
@@
We compared the performance among cTIRF, SOTA deconvolution (Sparse Deconvolution[zhao_sparse_2022]) and background removal (Dark Sectioning[cao_dark-based_2025]) methods, as shown in Fig.1 c. The Sparse Deconvolution gave a sharper reconstruction, while the out-of-focus signal still existed. The Dark Sectioning could efficiently remove the background, but barely enhanced the axial resolution, which was clearly shown by the region of the arch-like structure. It should also be noted that, ET2dNet was not specifically designed to implement deconvolution, but it was compatible with deconvolution methods, as shown in the sixth row in Fig.1 d. By directly implementing Sparse Deconvolution on the cTIRF, both lateral and axial resolution were enhanced. We further examined cTIRF on a test dataset containing 20 images, generating a better average MS-SSIM (compared to Sparse Deconvolution and Dark Sectioning), as shown in Fig.1 e. Additional examples of F-actin labeled U87 cells are shown in Extended Data Fig.1. We also compared other deconvolution and background removal methods, including RL deconvolution[richardson_bayesian-based_1972, lucy_iterative_1974], Wavelet-based background and noise subtraction (WBNS)[hupfel_wavelet-based_2021], Background Filtering (BF)[mo_quantitative_2023], while cTIRF outperformed all these methods, see Extended Data Fig.2. Brief introductions and implementation details of the above algorithms are provided in Supplementary Information Note 1. The details about the combination of cTIRF and Sparse deconvolution are provided in Supplementary Information Note 2.
The axial resolution enhancement achieved by ET2dNet was estimated using image decorrelation analysis [descloux_parameter-free_2019]. An EPI z-stack of an F-actin–labeled C6 cell was processed with ET2dNet (applied independently to each XY slice) to generate a cTIRF stack, and axial resolution was quantified on the corresponding XZ/YZ cross-sections using sectorial decorrelation analysis (Methods 4.7). This analysis indicates an average axial-resolution improvement of approximately threefold.
@@
The performance of cTIRF was further assessed using WF-confocal image pairs of human brain tissue sections stained with an antibody against GFAP (provided by ref.[wernersson_deconwolf_2024]), as shown in Fig.1 f. Note that the WF image was collected using 60X 1.4NA oil objective (with pixel size 107.9nm), which was slightly different from the training dataset using 60X 1.5NA oil objective (with pixel size 97.5nm). In such case, cTIRF could still efficiently remove background, outperforming the Sparse Deconvolution method and Dark Sectioning. As indicated by the region in the white box and the corresponding intensity profile, Sparse Deconvolution failed to suppress background, while Dark Sectioning erroneously eliminated signals. In contrast, cTIRF achieved a background-free result while preserving critical structural information, demonstrating high consistency with confocal GT.
2.2 Computational TIRF can be rapidly adapted to new imaging systems
@@
Our deep learning method aimed to generalize across WF microscopes featuring diverse objectives (e.g., varying numerical aperture and magnification). We hypothesized a pretrained ET2dNet exhibited robust performance under similar PSFs (though non-identical), as validated in Fig.1f. Here, input images acquired using a 60X/1.4NA objective were processed by a network trained on datasets acquired using a 60X/1.5NA objective. For systems exhibiting substantially different PSFs, fine-tuning was required. Implementation of the ET2dNet on WF images acquired with a 100X/1.5NA objective revealed effective background removal but negligible enhancement in axial-resolution, as shown in Fig.2a (denoted ET2dNet 60X). Specifically, at an arch-like structure (white dashed box), continuous microfilaments in the ET2dNet 60X output indicated a lack of axial-resolution improvement.
[…]
@@
To improve the axial-resolution enhancement capability of source ET2dNet, we finetuned it on a small dataset of 20 paired EPI-TIRF images. The finetuning process achieved rapid convergence, yielding ET2dNet 100X, which significantly outperformed ET2dNet 60X at axial-resolution enhancement. For instance, at the arch-like structure indicated by the white box in Fig.2 a, the ET2dNet 100X predicted discontinuous microfilaments, aligning with the measured TIRF. Quantitative validation across 20 test image pairs further confirmed improved PSNR and MS-SSIM after fine-tuning (Fig.2b).
@@
Furthermore, we demonstrated that the physics-informed hybrid network architecture enhanced generalization and advanced rapid adaptation. An ablation study investigating the impact of the self-supervised learning branch was presented in Fig.2 c and d. Both the baseline architecture model (without self-supervised learning branch) and the hybrid architecture model were trained on the 60X dataset and subsequently fine-tuned on the 100X dataset. Visually, the hybrid model regained its axial-resolution enhancement capability within 5 epochs during fine-tuning, whereas the baseline model required more than 20 epochs, as shown in Fig. 2c (the full images are provided in Supplementary Information Fig. S2). Statistical comparison in Fig. 2d revealed the hybrid model initially achieved higher average MS-SSIM (to experimental measured TIRF in the test dataset) and converged faster to a superior final value, reinforcing our conclusion. Ablation experiments conducted on networks with different model sizes (represented by S2, S1, S0 in Fig. 2d) consistently corroborated the above findings, underscoring the hybrid model’s strong transferability and superior generalization capacity.
@@
Axial resolution enhancement was further validated on images of stained human brain tissue sections (Fig. 2e). Comparative analysis of ET2dNet implementations (60X vs 100X) on z-stack (containing two 0.75 µm axial spacing slices) revealed critical performance differences. While the 60X model provided partial background suppression, the fine-tuned 100X model substantially eliminated out-of-focus signals. Notably, two microfilaments (indicated by green and blue arrows) located at the 0 µm and 0.75 µm planes exhibited cross-plane interference in EPI images. As evidenced by intensity profiles (Fig. 2e), both confocal microscopy and ET2dNet 60X failed to resolve these structures, whereas ET2dNet 100X successfully isolated the overlapping signals. This also demonstrated cTIRF can provide stronger optical sectioning than conventional confocal in our tested conditions.
2.3 Volumetric computational TIRF reconstruction